OpenStudy (anonymous):

How do you figure out the sin,cos,tan of equations such as 3 pi/4 without using calculator? I'm so confused. I know the answer for the sin of 3 pi/4 is negative root 2 over 2, but what are the steps to get those answers?

OpenStudy (campbell_st):

well you can put your calculator into radian mode rather than degree mode. or use the fact \[180^o = \pi\] so find \[\frac{3}{4} \times 180 = 135^o\] so this is a 2nd quadrant angle, where sine is positive. so the answer is \[\sin(\frac{3\pi}{4}) = \sin(135) = \sin(45) = \frac{\sqrt{2}}{2}\] hope it helps

OpenStudy (anonymous):

This clears it up a little bit and makes me understand it easier since my teacher just runs through problems instead of going step by step. Thank you so much!!!

OpenStudy (wolf1728):

There's a handy chart on this page: http://1728.org/degrees.htm that shows the PI value every 15° through 360°